OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6.

OpenStudy (anonymous):

put it in the form y = mx + b first

OpenStudy (anonymous):

Get in the form y=mx+c where m is the slope and c is the y-intercept: 3y=5x+6. Divide by 3: y=(5/3)x+2 So the slope is 5/3 and the y intercept is (0,2)

OpenStudy (anonymous):

Parallel lines have the same gradient, so change your equation to the form y=mx+c first y= -1/2 x + 1, therefore the slope is -1/2 13) 3. 14) 4. 15) 1. 16) 2. The rest of your questions uses up too much of my brain power I'm afraid xD

hero (hero):

Alternatively: Set x = 0 to get: 5(0) - 3y = -6 0 - 3y = -6 -3y = -6 6 = 3y 6/3 = y 2 = y